OpenStudy (anonymous):

Trig/ Pre Cal/ identities Am I on the right path? Please do not give answer but guide me. Thank you. cos(x - y) - cos(x+y) = 2sin x siny Is cos(x-y) like cos(a-B) = cos a cos B + sin a sin B

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cos(x-y) like cos(a-B) = cos a cos B + sin a sin B

OpenStudy (anonymous):

just plug x and y into a and b

OpenStudy (anonymous):

Ok, this does not seem to work. cos x *cos y + sin y *sin y - cos x * cos y - sin x * sin y = 2sin x 2 sin y Should be 2sin x sin y. Am I skipping something?

OpenStudy (anonymous):

cos(x+y) is not like cos(x-y)

OpenStudy (anonymous):

cos(x+y) = cos x cos y - sin x sin y

zepdrix (zepdrix):

yah your angle sum and different expansions look correct nixy :) do you see the next step after that?

zepdrix (zepdrix):

oh oh oh, i think you simply forgot to distribute the negative.

zepdrix (zepdrix):

\[\Large\rm \cos(x-y)\color{red}{-}\cos(x+y)\]\[\Large\rm =\cos x \cos y+ \sin x \sin y \color{red}{-}(\cos x \cos y -\sin x \sin y)\]

OpenStudy (anonymous):

Ah ok I see. man it was right in front of me. The way you put it made it clearer lol

OpenStudy (anonymous):

Thanks @zepdrix and @Thesn1

zepdrix (zepdrix):

np